12. Prove the following properties of Boolean algebras. Give a reason for each step. a. (x + y x) b. x.(z+y) + (x + y) = x
Answer #1

a)

(x+y.x)'

=x'.(y.x)' (demorgan's law)

=x'(y'+x') (demorgan's law)

=x'.y'+x'.x'(distributive law A.(B+C)=A.B+A.C)

=x'.y'+x' (Idempotent law A.A=A)

=x'+x'.y' (Commutative law)

=x' (absorption law A+AB=A where A=x', B=y')

b)

x.(z+y)+(x'+y)'

=x.z+x.y+(x'+y)' (distributive law)

=x.z+x.y+(x')'.(y') (demorgan's law)

=x.z+x.y+x.y' (double negation law (A')'=A)

=x.z+x.(y+y') (distributive law)

=x.z+x.1 (complement law A+A'=1)

=x.z+x

=x+xz

=x (absorption law)

c)

(x.y)'+x'.z+y'.z

=x'+y'+x'z+y'.z (demorgans law)

=x'+y'+z.(x'+y') (distributive law)

=x'+y'+(x'+y').z (commutative law)

=x'+y' (absorption law A+AB=A where A=x'+y' B=z)